OpenStudy (anonymous):

Does anyone know this? It says my calculations are off by 10% Vector B (with arrow above) has x, y, and z components of 6.00, 8.00, and 5.00 units, respectively. Calculate the magnitude of B (with arrow above). _11.18_ Calulcate the angle that makes with the x axis. ____° Calculate the angle that makes with the y axis. ____° Calculate the angle that makes with the z axis. ____°

OpenStudy (anonymous):

how much hw you got?

OpenStudy (anonymous):

last one lol

OpenStudy (anonymous):

but she gives us a lot with stuff we havent learned yet lol its why im on here

OpenStudy (anonymous):

11.1803

OpenStudy (anonymous):

i got that lol

OpenStudy (anonymous):

i need the rest

OpenStudy (phi):

or 5 sqrt(5)

OpenStudy (anonymous):

have you learned dot products yet?

OpenStudy (anonymous):

for angles i did arc tan of each of these 6/8 8/5 6/5 it told me i was wrong

OpenStudy (anonymous):

no

OpenStudy (phi):

acos

OpenStudy (phi):

e.g acos(6/11,18)

OpenStudy (anonymous):

{6,8,5} is your vector {1,0,0} is x axis \[\cos ^{-1}\left(\frac{A.B}{|A| |B|}\right)=\theta\] \[\text{ArcCos}\left[\frac{6}{5 \sqrt{5}}\right]\]=58 degree

OpenStudy (zarkon):

you can use the \vec command to make the arrow \[\vec{B}\]

OpenStudy (anonymous):

57.5437 to be more accurate

OpenStudy (anonymous):

ok so thats for x

OpenStudy (anonymous):

so for y is it arc sin?

OpenStudy (anonymous):

no, \[\left[\text{ArcCos}\left[\frac{8}{5 \sqrt{5}}\right]\right.\] =44.31

OpenStudy (anonymous):

so its a triangle right? so its 180-44.31-57.54?

OpenStudy (anonymous):

no, \[\text{ArcCos}\left[\frac{5}{5 \sqrt{5}}\right]\] for z which is 63.43

OpenStudy (anonymous):

i feel so dumb..i wish my teacher was better =( thanks for all your help today

OpenStudy (phi):

it is tough to do 3-dim. But if you focus on just one axis, say z. you have a line going straight up (z) and another slanting off (B) B is the hypotenuse

OpenStudy (anonymous):

did that check out?

OpenStudy (phi):

the cosine of the angle is z component/hypotenuse

OpenStudy (anonymous):

yea it all was correct! thanks!!! =)

OpenStudy (anonymous):

nice, so basically \[\text{ArcCos}\left[\frac{\text{Component}}{\text{Magnitude}}\right]\]